is v(v-3) = v^2 + 6 a quadratic equation and if so what is the solution?
Answer by kiru_khandelwal(79)

v(v-3) = v^2 + 6
=> v^2 - 3v = v^2 + 6
=> -3v = 6
=> 3v = -6
=> v = -2